# Optimal Deterministic Rendezvous in Labeled Lines

### v1

@@ Abstract
In a rendezvous task, a set of mobile agents initially dispersed in a network have to gather at an arbitrary common site. We consider the rendezvous problem on the infinite labeled line, with $2$ initially asleep agents, without communication, and a synchronous notion of time. Each node on the line is labeled with a unique positive integer. The initial distance between the two agents is denoted by $D$. Time is divided into rounds. We count time from the first moment that an agent wakes up, and denote by $\tau$ the delay in two agents’ wake up times. If awake in a given round $T$, an agent at a node $v$ has three options: stay at the node $v$, take port 00, or take port $1$. If it decides to stay, the agent will still be at node $v$ in round $T+1$. Otherwise, it will be at one of the two neighbors of $v$ on the infinite line, depending on the port it chose. The agents achieve rendezvous in $T$ rounds if they are at the same node in round $T$. We aim for a deterministic algorithm for this problem.
The problem was recently considered by Miller and Pelc [17, DISC 2023]. With $\ell_{\max}$ the largest label of the two starting nodes, they showed that no algorithm can guarantee rendezvous in $o(D\log^{*}\ell_{\max})$ rounds. The lower bound follows from a connection with the LOCAL model of distributed computing, and holds even if the agents are guaranteed simultaneous wake-up ($\tau=0$) and are told their initial distance $D$. Miller and Pelc also gave an algorithm of optimal matching complexity $O(D\log^{*}\ell_{\max})$ when the agents know $D$, but only obtained the higher bound of $O(D^{2}(\log^{*}\ell_{\max})^{3})$ when $D$ is unknown to the agents.
@@ 1 Introduction
A classic task for a set of autonomous mobile agents is to meet at some location. The agents might want to gather for a number of reasons: to exchange some resources, to initiate a long journey together, or to assemble in a specific shape. With only two agents, the problem is known as rendezvous [1].
We consider a fully discrete version of this problem: time is divided into rounds, and in each round, each agent is located at a node of a graph $G$. Furthermore, we assume that each node $v$ of this graph has a unique label $\ell(v)\in\mathbb{N}^{*}$, and that the graph $G$ inhabited by the agents is an infinite line. The agents wake up at possibly different rounds. Once awake, in each round, an agent can stay at its current location for one round or move to a node neighboring its current location. Rendezvous is achieved if the agents are at the same node in some round. In particular, it does not suffice that the agents are in adjacent nodes, or cross the same edge in opposite directions: they lack a communication or sensing method to detect these situations. The initial distance between the two agents is denoted $D$, and $\tau$ denotes the delay between the two wake-up times. The two agents execute the same algorithm, also known as the symmetric setting. We more formally describe our model in Section 2.
@@ 1.1 Our Results
Theorem 1.1 improves upon a prior algorithm of complexity $O(D^{2}(\log^{*}\ell_{\max})^{3})$ by Miller and and Pelc [17], where $\ell_{\max}$ is the largest label of the two starting nodes of the agents. To a lesser extent, we also improve upon their $O(D\log^{*}\ell_{\max})$ algorithm that assumes that the agents know their initial distance $D$, by replacing their dependency in $\ell_{\max}$ by one in $\ell_{\min}$.
The optimality of our algorithm follows from a lower bound from the same paper by Miller and Pelc [17]. They proved that any algorithm running in $o(D\log^{*}\ell_{\max})$ rounds would imply an $o(\log^{*}n)$ algorithm for $3$-coloring paths and cycles of $\operatorname{poly}(n)$ nodes in the LOCAL model. A seminal result by Linial states that no such $o(\log^{*}n)$ algorithm exists [14]. Note that our algorithm, while of complexity $o(D\log^{*}\ell_{\max})$ if considering an infinite family of instances where $\ell_{\min}\in o(\ell_{\max})$, is compatible with this lower bound, as families of instances where $\ell_{\min}\in\Theta(\ell_{\max})$ can be constructed. That is, our algorithm is only of complexity $o(D\log^{*}\ell_{\max})$ on specific instances, not in general [^1] Throughout this paper and as was the case in the paper by Miller and Pelc, we take a slightly non-standard definition of the $\log^{*}$ function guaranteeing that $\log^{*}(x)\geq 1$ for all $x$. This is to clarify that our $O(D\log^{*}\ell_{\min})$ algorithm does not finish in 00 rounds when $\ell_{\min}=1$, since $\log^{*}(1)=0$ with the most common definition of $\log^{*}$. See the Preliminaries (Section 2)..
@@
Like our algorithm for the infinite line, this improves upon previous results of complexity $O(\min(n,D^{2}(\log^{*}\ell_{\max})^{3}))$ and $O(\min(n,D\log^{*}\ell_{\max}))$ by Miller and and Pelc [17].
@@ 1.2 Technical Overview
Rendezvous and treasure hunt
@@
Coloring an infinite line
@@
Still, it is easy to adapt a seminal LOCAL algorithm for $3$-coloring paths and cycles to obtain a useful mapping. To do so, the algorithm is transformed into an early-stopping one, which has nodes terminate earlier depending on their given labels. While the classic LOCAL algorithm for $3$-coloring paths and cycles completes in $O(\log^{*}n)$ rounds, assuming that the nodes are given unique identifiers between $1$ and $\operatorname{poly}(n)$, the early-stopping version instead guarantees that a node $v$ of label $\ell(v)$ permanently commits to a color after $O(\log^{*}\ell(v))$ rounds. From the point of view of mapping local views to colors, this means that there exists a mapping that infers the color of a node of label $\ell(v)$ from its distance-$O(\log^{*}\ell(v))$-neighborhood. Used on the infinite sub-line obtained by taking one out of every $D$ nodes, this allows the two agents to assign distinct colors to their starting nodes after $O(D\log^{*}\ell_{\max})$ rounds of exploration, where $\ell_{\max}$ is the largest of the starting locations’ labels. This roughly summarizes the approach of Miller and Pelc [17].
Limitations of the regular sub-line approach
@@
As a result, to tackle the setting without knowledge of $D$ by using the previous algorithm with different guesses for $D$ basically requires to test every possible value for $D$. This is true even if the agents are given an upper bound on $D$. This difficulty is the source of the $O(D^{2})$ dependency in the earlier algorithm by Miller and Pelc [17].
Limitations of a purely coloring-based approach
@@
With only an upper bound on $D$, the agents now need to break symmetry with $2D$ nodes instead of just $2$. Guaranteeing that any two nodes within distance $D$ from one another are colored differently cannot be done with $3$ colors. Pushing forward with this approach, one could color the line with $O(D)$ colors. But now, the increased number of colors necessitates to design the same number of distinct behaviors of alternating searching and waiting phases guaranteeing rendezvous. This higher number of behaviors is likely to lead to a $D\log D$ dependency in the distance, similar to how in anonymous networks with labeled agents, larger labels lead to higher complexity in this fashion [^2] For instance, with $\ell$ the smallest of the two agents’ labels, $D$ their initial distance, and simultaneous startup, the complexity of rendezvous is $\Theta(D\log\ell)$ in the anonymous ring [7].. This would be far from the $\Omega(D\log^{*}\ell)$ lower bound.
Our solution: ruling set nodes as local landmarks
@@
A simplifying remark: avoiding simultaneous edge crossings
Contributing to the overall simplicity of our algorithm is a subroutine for crossing edges that prevents the agents from exchanging positions on the line without achieving rendezvous. The idea behind it is very simple: on the labeled line, every edge can be given a canonical orientation, for example, from lower to higher ID. By having the agents perform a different set of moves to cross an edge depending on the direction in which they take it, we can make sure that the agents must be at the same node at some point during the crossing. The first few moves are the same for both orientations of the edge, which enables an agent to discover the orientation of the edge while crossing it – e.g., on the labeled line, through learning the identifier of the other endpoint. Replacing every one-round edge crossing by this procedure only incurs a multiple factor of $4$ in the complexity of the algorithm. We believe that this observation might be useful to the design of future rendezvous algorithms on graphs whose edge can easily be given an orientation.
@@ 1.3 Related work
Most relevant to the results of this paper is the work by Miller and Pelc [17], that considered the rendezvous problem with the exact same set of assumptions as we do in this paper, as well as some easier settings. They showed that when agents know the distance $D$ between themselves, and the maximum labels of the two starting nodes of the agents is $\ell$, deterministic rendezvous can be done in $O(D\log^{*}\ell)$ rounds. They gave a matching lower bound of $\Omega(D\log^{*}\ell)$, shown on a line whose IDs all have the same $\log^{*}\ell$. The lower bound holds even with simultaneous startup. The upper bound does not assume this. When $D$ is unknown, they give an upper bound of $O(D^{2}(\log^{*}\ell)^{3})$. They also gave an algorithm to achieve rendezvous in $O(D)$ rounds when the agents know their positions on the infinite line, improving upon an earlier result by Collins et al. [5] which additionally assumed simultaneous startup of the agents.
On an anonymous line, where the agents themselves are given two distinct labels, the problem is known to be harder. Dessmark et al. [7] showed that the problem has complexity $\Theta(D\log\ell)$, assuming simultaneous startup, where $\ell$ is the smallest of the two agents’ labels. With arbitrary startup, they consider a harder setting where the agents only appear in the graph once they are awake. With that harder assumption, the algorithms tend to have a dependency on $n$, the size if the graph (which is then not assumed to be infinite in this work). Rendezvous in infinite lines was first studied in [16], in which the authors considered the harder asynchronous setting, but with the agents capable of detecting each other in edges.
{tblr}
[…]
@@
✓ ✓ ✓ ✓ ✓ ✓ ✓ ✓ $\Omega(D)$ folklore
✓ ✓ ✓ ✓ $\Theta(D\log\ell)$ [7]
(LoG) (LoG) ✓ ✓ ✓ $O(D)$ [5]
(LoG) (LoG) ✓ ✓ $O(D)$ [17]
✓ ✓ (NL) ✓ ✓ ✓ $\Omega(D\log^{*}\ell)$ [17]
✓ ✓ (NL) ✓ ✓ $\Theta(D\log^{*}\ell)$ [17]
✓ ✓ (NL) $O(D^{2}(\log^{*}\ell)^{3})$ [17]
✓ ✓ (NL) $\Theta(D\log^{*}\ell)$ this paper
Acronyms for assumptions: Lin (Graph is Line-like): the graph is $2$-regular, i.e., a line or a ring. $\mathclap{n{\ll}\infty}$ (Finite Graph): when unchecked, the graph can be infinite. NL (Node Labels): each node is equipped with a unique label. Implied by LoG, and implies AL. AL (Agent Labels): the agents are each given a unique label. Implied by LoG and NL. EwA (Exists while Asleep): rendezvous is also achieved if an awake agent finds an asleep agent. SWU (Simultaneous Wake-Up): the agents wake up in the same round ($\tau=0$). DbA (Distance between Agents): the agents know the initial distance $D$ between their starting locations. LoG (Location on Graph): the graph has some global naming scheme, which is known to the agents, and in which they know where they are. Implies NL and AL.
@@
Before they were formulated as questions of computational complexity and algorithm design, coordination problems of agents trying to meet appeared in the economics literature [20]. Algorithms for rendezvous in graphs were first studied by Alpern [2]. Some of the graph structures in which deterministic rendezvous has been considered include the infinite tree [4], the infinite grid [9], the plane [8] and the ring [11]. Deterministic rendezvous was also studied in general graphs [21], and in a setting where agents are given additional information about their location in the graph [5]. Rendezvous on the line has also been studied with random resources [13, 3].
The $\Omega(D\log^{*}\ell)$ lower bound obtained in [17] comes from a reduction of the rendezvous problem in labeled lines to the problem of coloring nodes on a line in the LOCAL model with some number of colors $C\geq 3$. This problem has been shown to have a $\Omega(\log^{*}n-\log^{*}C)$ lower bound [14, 15]. This lower bound is tight, due to an earlier algorithm by Cole and Vishkin matching its complexity [6]. Many works consider a version of the LOCAL model in which the nodes know an upper bound $n$ on the number of nodes, and have IDs between $1$ and $\operatorname{poly}(n)$. A generic way to remove this type of global knowledge was developed in a work by Korman, Sereni, and Viennot [12]. Several ideas are transferable from the setting where $n$ is unknown to that of infinite graphs.
Readers interested in a broader view of the area of deterministic rendezvous might be interested in a recent survey by Pelc [19]. The book by Alpern and Gal [1] offers an introduction to the area of rendezvous that includes many topics outside the scope of this paper (asymmetric agents, continuous versions of the problem). For an introduction to the LOCAL model, see the book by Peleg [18], as well as the online book by Hirvonen and Suomela [10].
@@ 1.4 Organization of the Paper
We start with some preliminaries in Section 2, where we introduce the notations and the model we are going to be considering throughout the paper.
In Section 3, we present the techniques and algorithms that the agents are going to be using to break symmetry between them and achieve rendezvous. Section 3.1 shows how to devise a way for the agents to meet during the execution of the algorithm when they both move through the same edge at the same time but from different starting nodes. In Section 3.2, we build an efficient algorithm to compute ruling sets on paths in LOCAL. Then, using this algorithm, we obtain in Section 3.3 an efficient algorithm for computing ruling sets on paths in LOCAL in which nodes of smaller IDs terminate earlier.
Section 4 describes an optimal algorithm for the agent to reach rendezvous in our setting and analyses its complexity. Section 4.1 presents the algorithm computed by the agents. To prove that the agent achieve rendezvous, we focus on a specific phase of the algorithm of one of the agents. We show that after reaching this phase the agent will achieve rendezvous before switching to the next phase. We split the analysis according to whether the agents are in the same phase or not. We argue about the two agents achieving rendezvous when in different phases in Section 4.2, and when in the same phase in Section 4.3. In Section 4.4, we show that the phase of our algorithm for which our arguments hold is guaranteed to occur within our stated upper bound, which finishes the proof of our optimal upper bound.

### v2

@@ Abstract
In a rendezvous task, a set of mobile agents initially dispersed in a network have to gather at an arbitrary common site. We consider the rendezvous problem on the infinite labeled line, with $2$ initially asleep agents, without communication, and a synchronous notion of time. Each node on the line is labeled with a unique positive integer. The initial distance between the two agents is denoted by $D$. Time is divided into rounds and measured from the moment an agent first wakes up. We denote by $\tau$ the delay between the two agents’ wake up times. If awake in a given round $T$, an agent at a node $v$ has three options: stay at the node $v$, take port 0, or take port $1$. If it decides to stay, the agent will still be at node $v$ in round $T+1$. Otherwise, it will be at one of the two neighbors of $v$ on the infinite line, depending on the port it chose. The agents achieve rendezvous in $T$ rounds if they are at the same node in round $T$. We aim for a deterministic algorithm for this problem.
The problem was recently considered by Miller and Pelc [MP25, Distributed Computing 2025]. With $\ell_{\max}$ the largest label of the two starting nodes, they showed that no algorithm can guarantee rendezvous in $o(D\log^{*}\ell_{\max})$ rounds. The lower bound follows from a connection with the LOCAL model of distributed computing, and holds even if the agents are guaranteed simultaneous wake-up ($\tau=0$) and are told their initial distance $D$. Miller and Pelc also gave an algorithm of optimal matching complexity $O(D\log^{*}\ell_{\max})$ when the agents know $D$, but only obtained the higher bound of $O(D^{2}(\log^{*}\ell_{\max})^{3})$ when $D$ is unknown to the agents.
@@ 1 Introduction
A classic task for a set of autonomous mobile agents is to meet at some location. The agents might want to gather for a number of reasons: to exchange some resources, to initiate a long journey together, or to assemble in a specific shape. With only two agents, the problem is known as rendezvous [AG03].
We consider a fully discrete version of this problem: time is divided into rounds, and in each round, each agent is located at a node of a graph $G$. Furthermore, we assume that each node $v$ of this graph has a unique label $\ell(v)\in\mathbb{N}^{*}$, and that the graph $G$ inhabited by the agents is an infinite line. The agents wake up at possibly different rounds. Once awake, in each round, an agent can stay at its current location for one round or move to a node neighboring its current location. Rendezvous is achieved if the agents are at the same node in some round. In particular, it does not suffice that the agents are in adjacent nodes, or cross the same edge in opposite directions: they lack a communication or sensing method to detect these situations. The initial distance between the two agents is denoted $D$, and $\tau$ denotes the delay between the two wake-up times. The two agents execute the same algorithm, also known as the symmetric setting. We more formally describe our model in Section˜2.
@@ 1.1 Our Results
Theorem˜1.1 improves upon a prior algorithm of complexity $O(D^{2}(\log^{*}\ell_{\max})^{3})$ by Miller and Pelc [MP25], where $\ell_{\max}$ is the largest label of the two starting nodes of the agents. To a lesser extent, we also improve upon their $O(D\log^{*}\ell_{\max})$ algorithm that assumes that the agents know their initial distance $D$, by replacing their dependency in $\ell_{\max}$ by one in $\ell_{\min}$.
The optimality of our algorithm follows from a lower bound from the same paper by Miller and Pelc [MP25]. They proved that any algorithm running in $o(D\log^{*}\ell_{\max})$ rounds would imply an $o(\log^{*}n)$ algorithm for $3$-coloring paths and cycles of $\operatorname{poly}(n)$ nodes in the LOCAL model. A seminal result by Linial states that no such $o(\log^{*}n)$ algorithm exists [Lin92]. Note that our algorithm, while of complexity $o(D\log^{*}\ell_{\max})$ if considering an infinite family of instances where $\ell_{\min}\in o(\ell_{\max})$, is compatible with this lower bound, as families of instances where $\ell_{\min}\in\Theta(\ell_{\max})$ can be constructed. That is, our algorithm is only of complexity $o(D\log^{*}\ell_{\max})$ on specific instances, not in general [^1] Throughout this paper and as was the case in the paper by Miller and Pelc, we take a slightly non-standard definition of the $\log^{*}$ function guaranteeing that $\log^{*}(x)\geq 1$ for all $x$. This is to clarify that our $O(D\log^{*}\ell_{\min})$ algorithm does not finish in 0 rounds when $\ell_{\min}=1$, since $\log^{*}(1)=0$ with the most common definition of $\log^{*}$. See the Preliminaries (Section 2)..
@@
Like our algorithm for the infinite line, this improves upon previous results of complexity $O(\min(n,D^{2}(\log^{*}\ell_{\max})^{3}))$ and $O(\min(n,D\log^{*}\ell_{\max}))$ by Miller and Pelc [MP25].
@@ 1.2 Technical Overview
Rendezvous and treasure hunt.
@@
Coloring an infinite line.
@@
Still, it is easy to adapt a seminal LOCAL algorithm for $3$-coloring paths and cycles to obtain a useful mapping. To do so, the algorithm is transformed into an early-stopping one, which has nodes terminate earlier depending on their given labels. While the classic LOCAL algorithm for $3$-coloring paths and cycles completes in $O(\log^{*}n)$ rounds, assuming that the nodes are given unique identifiers between $1$ and $\operatorname{poly}(n)$, the early-stopping version instead guarantees that a node $v$ of label $\ell(v)$ permanently commits to a color after $O(\log^{*}\ell(v))$ rounds. From the point of view of mapping local views to colors, this means that there exists a mapping that infers the color of a node of label $\ell(v)$ from its distance-$O(\log^{*}\ell(v))$-neighborhood. Used on the infinite sub-line obtained by taking one out of every $D$ nodes, this allows the two agents to assign distinct colors to their starting nodes after $O(D\log^{*}\ell_{\max})$ rounds of exploration, where $\ell_{\max}$ is the largest of the starting locations’ labels. This roughly summarizes the approach of Miller and Pelc [MP25].
Limitations of the regular sub-line approach.
@@
As a result, to tackle the setting without knowledge of $D$ by using the previous algorithm with different guesses for $D$ basically requires to test every possible value for $D$. This is true even if the agents are given an upper bound on $D$. This difficulty is the source of the $O(D^{2})$ dependency in the earlier algorithm by Miller and Pelc [MP25].
Limitations of a purely coloring-based approach.
@@
With only an upper bound on $D$, the agents now need to break symmetry with $2D$ nodes instead of just $2$. Guaranteeing that any two nodes within distance $D$ from one another are colored differently cannot be done with $3$ colors. Pushing forward with this approach, one could color the line with $O(D)$ colors. But now, the increased number of colors necessitates to design the same number of distinct behaviors of alternating searching and waiting phases guaranteeing rendezvous. This higher number of behaviors is likely to lead to a $D\log D$ dependency in the distance, similar to how in anonymous networks with labeled agents, larger labels lead to higher complexity in this fashion [^2] For instance, with $\ell$ the smallest of the two agents’ labels, $D$ their initial distance, and simultaneous startup, the complexity of rendezvous is $\Theta(D\log\ell)$ in the anonymous ring [DFKP06].. This would be far from the $\Omega(D\log^{*}\ell)$ lower bound.
Our solution: ruling set nodes as local landmarks.
@@
A simplifying remark: avoiding simultaneous edge crossings.
Contributing to the overall simplicity of our algorithm is a subroutine for crossing edges that prevents the agents from exchanging positions on the line without achieving rendezvous. The idea behind it is very simple: on the labeled line, every edge can be given a canonical orientation, for example, from lower to higher ID. By having the agents perform a different set of moves to cross an edge depending on the direction in which they take it, we can make sure that the agents must be at the same node at some point during the crossing. The first few moves are the same for both orientations of the edge, which enables an agent to discover the orientation of the edge while crossing it – e.g., on the labeled line, through learning the identifier of the other endpoint. Replacing every one-round edge crossing by this procedure only incurs a multiplicative factor of $4$ in the complexity of the algorithm. We believe that this observation might be useful to the design of future rendezvous algorithms on graphs whose edges can easily be given an orientation.
@@ 1.3 Related Work
Most relevant to the results of this paper is the work by Miller and Pelc [MP25], that considered the rendezvous problem with the exact same set of assumptions as we do in this paper, as well as some easier settings. They showed that when agents know the distance $D$ between themselves, and the maximum labels of the two starting nodes of the agents is $\ell$, deterministic rendezvous can be done in $O(D\log^{*}\ell)$ rounds. They gave a matching lower bound of $\Omega(D\log^{*}\ell)$, shown on a line whose IDs all have the same $\log^{*}\ell$. The lower bound holds even with simultaneous startup. The upper bound does not assume this. When $D$ is unknown, they give an upper bound of $O(D^{2}(\log^{*}\ell)^{3})$. They also gave an algorithm to achieve rendezvous in $O(D)$ rounds when the agents know their positions on the infinite line, improving upon an earlier result by Collins et al. [CCG+11] which additionally assumed simultaneous startup of the agents.
On an anonymous line, where the agents themselves are given two distinct labels, the problem is known to be harder. Dessmark et al. [DFKP06] showed that the problem has complexity $\Theta(D\log\ell)$, assuming simultaneous startup, where $\ell$ is the smallest of the two agents’ labels. With arbitrary startup, they consider a harder setting where the agents only appear in the graph once they are awake. With that harder assumption, the algorithms tend to have a dependency on $n$, the size if the graph (which is then not assumed to be infinite in this work). Rendezvous in infinite lines was first studied in [MGK+06], in which the authors considered the harder asynchronous setting, but with the agents capable of detecting each other in edges.
Assumptions Bound Citation
@@
$\Omega(D)$ folklore
$\Theta(D\log\ell)$ [DFKP06]
$O(D)$ [CCG+11]
$O(D)$ [MP25]
$\Omega(D\log^{*}\ell)$ [MP25]
$\Theta(D\log^{*}\ell)$ [MP25]
$O(D^{2}(\log^{*}\ell)^{3})$ [MP25]
$\Theta(D\log^{*}\ell)$ this paper
Acronyms for assumptions: Lin (Graph is Line-like): the graph has maximum degree $2$. $n{\ll}\infty$ (Finite Graph): when unchecked, the graph can be infinite. NL (Node Labels): each node is equipped with a unique label. Implied by LoG, and implies AL. AL (Agent Labels): the agents are each given a unique label. Implied by LoG and NL. EwA (Exists while Asleep): rendezvous is also achieved if an awake agent finds an asleep agent. SWU (Simultaneous Wake-Up): the agents wake up in the same round ($\tau=0$). DbA (Distance between Agents): the agents know the initial distance $D$ between their starting locations. LoG (Location on Graph): the graph has some global naming scheme, which is known to the agents, and in which they know where they are. Implies NL and AL.
@@
Before they were formulated as questions of computational complexity and algorithm design, coordination problems of agents trying to meet appeared in the economics literature [Sch81]. Algorithms for rendezvous in graphs were first studied by Alpern [Alp95]. Some of the graph structures in which deterministic rendezvous has been considered include the infinite tree [BP24], the infinite grid [GP25], the plane [DPP23] and the ring [KSSK03]. Deterministic rendezvous was also studied in general graphs [TZ14], and in a setting where agents are given additional information about their location in the graph [CCG+11]. Rendezvous on the line has also been studied with random resources [LA96, BG98].
The $\Omega(D\log^{*}\ell)$ lower bound obtained in [MP25] comes from a reduction of the rendezvous problem in labeled lines to the problem of coloring nodes on a line in the LOCAL model with some number of colors $C\geq 3$. This problem has been shown to have a $\Omega(\log^{*}n-\log^{*}C)$ lower bound [Lin92, LS14]. This lower bound is tight, due to an earlier algorithm by Cole and Vishkin matching its complexity [CV86]. Many works consider a version of the LOCAL model in which the nodes know an upper bound $n$ on the number of nodes, and have IDs between $1$ and $\operatorname{poly}(n)$. A generic way to remove this type of global knowledge was developed in a work by Korman, Sereni, and Viennot [KSV13]. Several ideas are transferable from the setting where $n$ is unknown to that of infinite graphs.
Readers interested in a broader view of the area of deterministic rendezvous might be interested in a recent survey by Pelc [Pel19]. The book by Alpern and Gal [AG03] offers an introduction to the area of rendezvous that includes many topics outside the scope of this paper (asymmetric agents, continuous versions of the problem). For an introduction to the LOCAL model, see the book by Peleg [Pel00], as well as the online book by Hirvonen and Suomela [HS20].
@@ 1.4 Organization of the Paper
We start with some preliminaries in Section˜2, where we introduce the notations and the model we are going to be considering throughout the paper.
In Section˜3, we present the techniques and algorithms that the agents are going to be using to break symmetry between them and achieve rendezvous. Section˜3.1 shows how to devise a way for the agents to meet during the execution of the algorithm when they both move through the same edge at the same time but from different starting nodes. In Section˜3.2, we build an efficient algorithm to compute ruling sets on paths in LOCAL. Then, using this algorithm, we obtain in Section˜3.3 an efficient algorithm for computing ruling sets on paths in LOCAL in which nodes with smaller IDs terminate earlier.
Section˜4 describes an optimal algorithm for the agent to reach rendezvous in our setting and analyses its complexity. Section˜4.1 presents the algorithm computed by the agents. To prove that the agents achieve rendezvous, we focus on a specific phase of the algorithm of one of the agents. We show that after reaching this phase the agents will achieve rendezvous before switching to the next phase. We split the analysis according to whether the agents are in the same phase or not. We argue about the two agents achieving rendezvous when in different phases in Section˜4.3, and when in the same phase in Section˜4.4. We additionally show that the phase of our algorithm for which our arguments hold is guaranteed to occur within our stated upper bound, and finish the proof of our optimal upper bound in Section˜4.5.
We give in Section˜5 the implications of our results for line-like graphs other than the infinite line: finite lines, finite cycles, and the infinite half-line (containing a single node of degree $1$).